can you please help me write the equation in standard form for the parabola with vertex (0,0) and the directrix y=-11
Answer by ewatrrr(24785)


Hi,
Note: for a parabola, 4p(y-k) = (x-h)^2 where (h,k) is the vertex and
p the distance from the focus to the vertex. in this ex vertex is Pt(0,0)
Directrix y = -11 tells us p = 11 as the directrix is the same distance
from the vertex as is the focus.
4p(y-k) = (x-h)^2
4*11y = x^2
y = x^2/44 the standard form for this parabola